OpenStudy (anonymous):

Which of the following is the point slope equation of a line with a slope = 4/5 and a point of (-2, 1) ?

OpenStudy (anonymous):

@Chelsea04

OpenStudy (anonymous):

didn't I just answer this question?

OpenStudy (anonymous):

I got B :)

OpenStudy (anonymous):

oops! soooo sorry hold on

OpenStudy (anonymous):

in any case, find the slope as stated: 4/5 then substitute it into the equation: y-y1=m(x-x1) the point: (-2,1) -2=x1 1=y1

OpenStudy (anonymous):

here I'll post the right one

OpenStudy (anonymous):

get this equation into the form y=mx+b

OpenStudy (anonymous):

y-3=9(x+10) y=9x+90+3

OpenStudy (anonymous):

@Chelsea04 so that's the form?

OpenStudy (anonymous):

do you gave to solve?

OpenStudy (anonymous):

not quite, from y=9x+90+3 you simplify to y=9x+93, then with the equation define m and b

OpenStudy (anonymous):

m is 9 b is 93

OpenStudy (anonymous):

oohhhh okay! So you add m and B? :)

OpenStudy (anonymous):

no, no need

OpenStudy (anonymous):

m remember is the slope b is the y intercept

OpenStudy (anonymous):

you just need to take it from the equation, so the answer is 9

OpenStudy (anonymous):

Ohh thank you sooo much! I'm almost finished <3 <3

OpenStudy (anonymous):

next?